OpenStudy (anonymous):

Find the distance between the points (6, 5) and (4, 3).

OpenStudy (mathstudent55):

Are you familiar with the distance formula?

OpenStudy (anonymous):

\[\large Distance = \sqrt{(x_2 - x_1)^2 + (y_2 - y_1)^2}\]

OpenStudy (anonymous):

Here : x1 = 6, x2 = 4, y1 = 5 and y2 = 3.

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

Just implement that formula and post your answer here so that we can check whether you got the right answer..

OpenStudy (anonymous):

2?

OpenStudy (anonymous):

No.. Check it once more..

OpenStudy (anonymous):

idk ? :,(

OpenStudy (anonymous):

See: \[D = \sqrt{(4-6)^2 + (3 - 5)^2} \implies \sqrt{(-2)^2 + (-2)^2}\]

OpenStudy (anonymous):

Can you tell me : \[(-2)^2 = ??\]

OpenStudy (anonymous):

4?

OpenStudy (anonymous):

Yep: But there are two (-2)^2 there: So you will get two 4..

OpenStudy (anonymous):

\[\sqrt{4 + 4} = ??\]

OpenStudy (anonymous):

Solve this on your own now..